mathslover (mathslover):

I want to know that can i prove heron's formula by applying pythagoras theorem

mathslover (mathslover):

@vishweshshrimali5

OpenStudy (vishweshshrimali5):

\begin{array}l\color{red}{\normalsize\text{J}}\color{orange}{\normalsize\text{u}}\color{#9c9a2e}{\normalsize\text{s}}\color{green}{\normalsize\text{t}}\color{blue}{\normalsize\text{ }}\color{purple}{\normalsize\text{w}}\color{purple}{\normalsize\text{a}}\color{red}{\normalsize\text{i}}\color{orange}{\normalsize\text{t}}\color{#9c9a2e}{\normalsize\text{ }}\color{green}{\normalsize\text{l}}\color{blue}{\normalsize\text{e}}\color{purple}{\normalsize\text{t}}\color{purple}{\normalsize\text{ }}\color{red}{\normalsize\text{m}}\color{orange}{\normalsize\text{e}}\color{#9c9a2e}{\normalsize\text{ }}\color{green}{\normalsize\text{t}}\color{blue}{\normalsize\text{h}}\color{purple}{\normalsize\text{i}}\color{purple}{\normalsize\text{n}}\color{red}{\normalsize\text{k}}\color{orange}{\normalsize\text{ }}\color{#9c9a2e}{\normalsize\text{a}}\color{green}{\normalsize\text{b}}\color{blue}{\normalsize\text{o}}\color{purple}{\normalsize\text{u}}\color{purple}{\normalsize\text{t}}\color{red}{\normalsize\text{ }}\color{orange}{\normalsize\text{t}}\color{#9c9a2e}{\normalsize\text{h}}\color{green}{\normalsize\text{i}}\color{blue}{\normalsize\text{s}}\color{purple}{\normalsize\text{}}\end{array}

OpenStudy (anonymous):

yes, you can.

mathslover (mathslover):

But how ?

OpenStudy (anonymous):

The question only asks if you can or not prove heron's formula for area using Pythagoras theorem. It doesn't asks how...

mathslover (mathslover):

:D so u not know ?

OpenStudy (foolaroundmath):

 \(CD = \sqrt{b^{2}-h^{2}} , BD = \sqrt{c^{2}-h^{2}}, a = CD+BD,Area= ah/2\) Using these 4 statements, we can prove heron's formula. (additional info required: area = 0.5*base*height) Hint: try finding \(h\) in terms of \(a,b,c\) from first three relations then substitute that in the area formula